# FIARSE: Model-Heterogeneous Federated Learning via Importance-Aware Submodel Extraction

### v1

@@ Appendix A Related Work
In addition to submodel extraction, another effective method for facilitating model heterogeneity is model customization [38, 60, 21, 9, 53, 62]. This approach allows clients to construct local models that align with their local computation resources. Since the clients’ local models could be very distinct, the existing works implement model aggregation through the technique of knowledge distillation (KD) [14]. In detail, each client first locally trains its model on a shared public dataset, and then sends the model’s output (i.e., logits) to the server. The server gathers and aggregates clients’ outputs into a unified model output and broadcasts it to the clients [49]. Prior to the next round of local training, the clients thereby fine-tune the local models by incorporating the aggregated logits. Although the process effectively transfers knowledge across heterogeneous models, model customization requires a shared public dataset that should have a similar distribution with the clients’ training data, which is unattainable in some cases [48, 3]. In our work, we do not require a public dataset, rendering it applicable to a broader range of applications.
@@
Up to now, Flado [37] is the first and only work to realize that the server tailors the submodels to align with the clients’ capacities. It achieves the performance trade-off between the global model and the clients’ submodels by differentiating model parameters based on their importance levels. However, this advantage comes at the cost of requiring clients to explicitly maintain the importance score of each model parameter. This requirement, consequently, introduces a considerable overhead on each client in terms of both storage and computation, which imposes great challenges on resource-constrained devices. Contrasting with Flado, our FIARSE explores the correlation between a model parameter’s importance level and its value. It implicitly represents the importance level through the parameter’s value, thereby eliminating the need to explicitly maintain separate importance scores for each model parameter.
@@ Appendix B Derivation of Equation (3)
$\frac{\partial\mathcal{M}^{(i)}}{\partial\tilde{x}}$ $=\cup_{j}\frac{\partial}{\partial\tilde{x}_{j}}(\frac{|\tilde{x}_{j}|-\theta_{i}}{|\tilde{x}_{j}|+\theta_{i}})\cdot\boldsymbol{1}_{|\tilde{x}_{j}|\geq\theta_{i}}$ (6)
$=\cup_{j}\frac{\partial|\tilde{x}_{j}|}{\partial x_{j}}\cdot\frac{\partial}{\partial|\tilde{x}_{j}|}(\frac{|\tilde{x}_{j}|-\theta_{i}}{|\tilde{x}_{j}|+\theta_{i}})\cdot\boldsymbol{1}_{|\tilde{x}_{j}|\geq\theta_{i}}$ (7)
$=\cup_{j}\frac{\partial|\tilde{x}_{j}|}{\partial x_{j}}\cdot\frac{2\theta_{i}}{(|\tilde{x}_{j}|+\theta_{i})^{2}}\cdot\boldsymbol{1}_{|\tilde{x}_{j}|\geq\theta_{i}}$ (8)
@@
$=\frac{\partial F_{i}(\tilde{x}\odot\mathcal{M}^{(i)})}{\partial\tilde{x}\odot\mathcal{M}^{(i)}}\odot(\mathcal{M}^{(i)}+\tilde{x}\odot\frac{\partial\mathcal{M}^{(i)}}{\partial\tilde{x}})$ (10)
$=\nabla F_{i}(\tilde{x}\odot\mathcal{M}^{(i)})\odot\mathcal{M}^{(i)}(1+\frac{2|\tilde{x}|\theta_{i}}{(|\tilde{x}|+\theta_{i})^{2}})$ (11)
@@ C.1 Useful Lemmas
[Lemma 1 in [56]] For a vector $v\in\mathbb{R}^{d}$, we partition it into $n$ sets, where we present it in the following table:
@@ C.2 Preliminary
Suppose that Assumption 4.1, 4.2 and 4.3 hold. Let the local learning rate satisfy $\eta_{l}\leq\frac{1}{2L\sqrt{K(K+1)}}$. With FIARSE, we have the following conclusion:
@@
First, we try to build a recursive relation between $\mathbb{E}\left\|\boldsymbol{x}_{t,k+1}^{(i)}-\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}$ and $\mathbb{E}\left\|\boldsymbol{x}_{t,k}^{(i)}-\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}$:
@@
$\quad+4(1+K)\eta_{l}^{2}\left(L^{2}\mathbb{E}\left\|\boldsymbol{x}_{t,k}^{(i)}-\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}+\delta^{2}\left\|\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}\right)$
$\quad+4(1+K)\eta_{l}^{2}\left(\mathbb{E}\left\|\nabla F_{i}^{[0:\gamma_{i}]}\left(\tilde{\boldsymbol{x}}_{t}\right)-\nabla F^{[0:\gamma_{i}]}\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}+\mathbb{E}\left\|\nabla F^{[0:\gamma_{i}]}\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}\right)$ (24)
where Equation (17) can be expanded to four terms, i.e., Equation (19), Equation (20), Equation (21), and Equation (22). Furthermore, Equation (24) is obtained by applying Assumption 4.1 to Equation (19), Assumption 4.3 to Equation (20), and two facts that $\nabla_{\tilde{\boldsymbol{x}}_{t}}F_{i}\left(\tilde{\boldsymbol{x}}_{t}\right)\odot\mathcal{M}^{(i)}_{t}\left(\tilde{\boldsymbol{x}}_{t}\right)=\nabla F_{i}^{[0:\gamma_{i}]}\left(\tilde{\boldsymbol{x}}_{t}\right)$ and $\nabla F\left(\tilde{\boldsymbol{x}}_{t}\right)\odot\mathcal{M}^{(i)}_{t}\left(\tilde{\boldsymbol{x}}_{t}\right)=\nabla F^{[0:\gamma_{i}]}\left(\tilde{\boldsymbol{x}}_{t}\right)$ to Equation (21) and Equation (22).
As $\eta_{l}^{2}\leq\frac{1}{2L\sqrt{K(K+1)}}$, we have
@@
where
$C=\delta^{2}\left\|\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}+\mathbb{E}\left\|\nabla F_{i}^{[0:\gamma_{i}]}\left(\tilde{\boldsymbol{x}}_{t}\right)-\nabla F^{[0:\gamma_{i}]}\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}+\mathbb{E}\left\|\nabla F^{[0:\gamma_{i}]}\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}.$ (28)
Equation (25) establishes because $\left(1+\frac{2}{K}\right)^{k}\leq 9$ for all $k\in\mathbb{N}_{+}$. Therefore, for all $k\in[K]$, we have
@@
With the above conclusion, we can find the bound of LHS of Equation (15), which is
@@
By applying Assumption 4.2 to Equation (32), we can attain the desired conclusion. ∎
@@
Suppose that Assumption 4.1, 4.2 and 4.3 hold. Let the local learning rate satisfy $\eta_{l}\leq\min\left(\frac{1}{2L\sqrt{K(K+1)}},\frac{1}{6L\sqrt{(K+1)A}}\right)$. With FIARSE, we have the following conclusion:
$\quad\sum_{j=0}^{n}\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\cdot\mathbb{E}\left\|\nabla F^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right)-\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|K}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right)\right\|_{2}^{2}$ (33)
$\leq 4A\delta^{2}\left\|\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}+72KA\eta_{l}^{2}L^{2}\left(\sigma^{2}+\left\|\nabla F\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}\right)$ (34)
@@
First of all, let us find the bound for a single term of Equation (33):
$\quad\mathbb{E}\left\|\nabla F^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right)-\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|K}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right)\right\|_{2}^{2}$ (35)
@@
$\leq 2\mathbb{E}\left\|\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|}\sum_{i\in N_{\gamma_{j}^{\prime}}}\left(\nabla F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right)-\nabla_{\tilde{\boldsymbol{x}}_{t}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\odot\mathcal{M}^{(i)}_{t}\left(\tilde{\boldsymbol{x}}_{t}\right)\right)\right)\right\|_{2}^{2}$ (36)
$\quad+2\mathbb{E}\left\|\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|K}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\left(\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right)-\nabla_{\tilde{\boldsymbol{x}}_{t}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\odot\mathcal{M}^{(i)}_{t}\left(\tilde{\boldsymbol{x}}_{t}\right)\right)\right)\right\|_{2}^{2}$ (37)
@@
$\leq 2\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|}\sum_{i\in N_{\gamma_{j}^{\prime}}}\mathbb{E}\left\|\nabla F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right)-\nabla_{\tilde{\boldsymbol{x}}_{t}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\odot\mathcal{M}^{(i)}_{t}\left(\tilde{\boldsymbol{x}}_{t}\right)\right)\right\|_{2}^{2}$ (38)
$\quad+2\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|K}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\mathbb{E}\left\|\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right)-\nabla_{\tilde{\boldsymbol{x}}_{t}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\odot\mathcal{M}^{(i)}_{t}\left(\tilde{\boldsymbol{x}}_{t}\right)\right)\right\|_{2}^{2}$ (39)
The last inequality is built upon Cauchy–Schwarz inequality. It is easily to have the following conclusion:
$\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|}\cdot\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\leq\frac{A}{N};\quad\frac{A}{N}\leq\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\leq 1$ (40)
Next, we sum Equation (35) for all $j$ through its segment and attain:
$\quad\sum_{j=0}^{n}\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\cdot\mathbb{E}\left\|\nabla F^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right)-\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|K}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right)\right\|_{2}^{2}$ (41)
$\leq 2\sum_{j=0}^{n}\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}\left|N_{\gamma_{j}^{\prime}}\right|}\sum_{i\in N_{\gamma_{j}^{\prime}}}\mathbb{E}\left\|\nabla F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right)-\nabla_{\tilde{\boldsymbol{x}}_{t}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\odot\mathcal{M}^{(i)}_{t}\left(\tilde{\boldsymbol{x}}_{t}\right)\right)\right\|_{2}^{2}$ (42)
$\quad+2\sum_{j=0}^{n}\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}\left|N_{\gamma_{j}^{\prime}}\right|K}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\mathbb{E}\left\|\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right)-\nabla_{\tilde{\boldsymbol{x}}_{t}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\odot\mathcal{M}^{(i)}_{t}\left(\tilde{\boldsymbol{x}}_{t}\right)\right)\right\|_{2}^{2}$ (43)

### v2

@@ Appendix A Related Work
Computation Heterogeneity in FL.
Computation heterogeneity in FL has become a critical area of research due to the varying computational capacities of participating clients. In traditional FL setups, each client is expected to perform the same volume of computation, such as training on local datasets for a fixed number of iterations, regardless of their hardware capabilities or resource availability [14, 15, 18, 63, 24, 68]. This assumption leads to inefficiencies, particularly when slower devices cause stragglers that delay model aggregation [44, 64, 23]. Recent work has proposed adaptive aggregation strategies to handle computation heterogeneity by assigning different workloads or varying the number of local updates based on client resources, allowing faster clients to contribute more to the global model while slower clients contribute less frequently [64, 43, 53, 50, 57, 72]. However, these works require the clients to train the full model, so they are infeasible when some of the clients cannot load the full model due to limited computation resources. In contrast, our work extracts a submodel for each client that fits within their computational capacity.
@@
In addition to submodel extraction, another effective method for facilitating model heterogeneity is model customization [46, 77, 28, 11, 66, 80]. This approach allows clients to construct local models that align with their local computation resources. Since the clients’ local models could be very distinct, the existing works implement model aggregation through the technique of knowledge distillation [19, 61, 46, 78]. In detail, each client first locally trains its model on a shared public dataset, and then sends the model’s output (i.e., logits) to the server. The server gathers and aggregates clients’ outputs into a unified model output and broadcasts it to the clients [60]. Prior to the next round of local training, the clients thereby fine-tune the local models by incorporating the aggregated logits. Although the process effectively transfers knowledge across heterogeneous models, model customization requires a shared public dataset that should have a similar distribution with the clients’ training data, which is unattainable in some cases [59, 3]. In our work, we do not require a public dataset, rendering it applicable to a broader range of applications.
@@
An existing work, Flado [45], realizes that the server tailors the submodels to align with various computation resources among clients. It achieves the performance trade-off between the global model and the clients’ submodels by differentiating model parameters based on their importance levels. However, this advantage comes at the cost of requiring clients to explicitly maintain the importance score of each model parameter. This requirement, consequently, introduces a considerable overhead on each client in terms of both storage and computation, which imposes great challenges on resource-constrained devices. Contrasting with Flado, our FIARSE explores the correlation between a model parameter’s importance level and its value. It implicitly represents the importance level through the parameter’s value, thereby eliminating the need to explicitly maintain separate importance scores for each model parameter.
@@ Appendix B Derivation of Equation (3)
$\frac{\partial\mathcal{M}^{(i)}}{\partial\tilde{x}}$ $=\bigcup_{j}\frac{\partial}{\partial\tilde{x}_{j}}\left(\frac{|\tilde{x}_{j}|-\theta_{i}}{|\tilde{x}_{j}|+\theta_{i}}\right)\cdot\boldsymbol{1}_{|\tilde{x}_{j}|\geq\theta_{i}}$ (6)
$=\bigcup_{j}\frac{\partial|\tilde{x}_{j}|}{\partial x_{j}}\cdot\frac{\partial}{\partial|\tilde{x}_{j}|}\left(\frac{|\tilde{x}_{j}|-\theta_{i}}{|\tilde{x}_{j}|+\theta_{i}}\right)\cdot\boldsymbol{1}_{|\tilde{x}_{j}|\geq\theta_{i}}$ (7)
$=\bigcup_{j}\frac{\partial|\tilde{x}_{j}|}{\partial x_{j}}\cdot\frac{2\theta_{i}}{(|\tilde{x}_{j}|+\theta_{i})^{2}}\cdot\boldsymbol{1}_{|\tilde{x}_{j}|\geq\theta_{i}}$ (8)
@@
$=\frac{\partial F_{i}(\tilde{x}\odot\mathcal{M}^{(i)})}{\partial\tilde{x}\odot\mathcal{M}^{(i)}}\odot\left(\mathcal{M}^{(i)}+\tilde{x}\odot\frac{\partial\mathcal{M}^{(i)}}{\partial\tilde{x}}\right)$ (10)
$=\nabla F_{i}(\tilde{x}\odot\mathcal{M}^{(i)})\odot\mathcal{M}^{(i)}\left(1+\frac{2|\tilde{x}|\theta_{i}}{(|\tilde{x}|+\theta_{i})^{2}}\right)$ (11)
@@ C.1 Useful Lemmas
For a vector $v\in\mathbb{R}^{d}$, we partition it into $n$ sets, where we present it in Table 2.
@@
Proof.
For a given submodel $v_{\left[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}\right]}$, there are $\left|N_{\gamma_{j}^{\prime}}\right|$ clients holding the submodel. In other words, $\left(N-\left|N_{\gamma_{j}^{\prime}}\right|\right)$ clients do not hold any parameters of $v_{\left[\gamma_{j-1}^{\prime}:1\right]}$. If a selected client $i\in\mathcal{A}$ holds the submodel $v_{\left[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}\right]}$, it will be divided by an integer $k\in\{1,\dots,A\}$ (i.e., $\left|\mathcal{A}_{\gamma_{j}^{\prime}}\right|=k$) with a probability of $\frac{\binom{\left|N_{\gamma_{j}^{\prime}}\right|-1}{k-1}\cdot\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A-k}}{\binom{N}{A}}$. Therefore, this part is expected to obtain the following result after aggregation:
$\mathbb{E}\textsf{Agg}_{i\in\mathcal{A}}\left(v_{\left[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}\right]}\right)$ $=\sum_{k=1}^{A}\frac{\binom{\left|N_{\gamma_{j}^{\prime}}\right|-1}{k-1}\cdot\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A-k}}{\binom{N}{A}}\cdot\frac{1}{k}\sum_{i\in N_{\gamma_{j}^{\prime}}}v_{\left[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}\right]}^{(i)}$ (15)
$=\sum_{k=1}^{A}\frac{\binom{\left|N_{\gamma_{j}^{\prime}}\right|}{k}\cdot\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A-k}}{\binom{N}{A}}\cdot\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|}\sum_{i\in N_{\gamma_{j}^{\prime}}}v_{\left[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}\right]}^{(i)}$ (16)
$=\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\cdot\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|}\sum_{i\in N_{\gamma_{j}^{\prime}}}v_{\left[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}\right]}^{(i)}$ (17)
By concatenating all $j\in[n]$, we can attain the desired results. ∎
In addition to the above lemma, we provide a bound for $\frac{\binom{N}{A}-\binom{N-c}{A}}{\binom{N}{A}}$ and $\frac{\binom{N}{A}-\binom{N-c}{A}}{c\cdot\binom{N}{A}}$, where $c$ is an integer of $\{1,\dots,N-A\}$.
Lemma C.2.
Given $N$ and $A$ are two integers where $A\leq N$, and $c$ is an integer of $\{1,\dots,N-A\}$, the following two inequalities are always true:
$\frac{A}{N}\leq\frac{\binom{N}{A}-\binom{N-c}{A}}{\binom{N}{A}}\leq 1;\qquad\frac{\binom{N}{A}-\binom{N-c}{A}}{c\cdot\binom{N}{A}}\leq\frac{A}{N}$ (18)
Proof.
For the first inequality, the lower bound holds when $c=1$, while the upper bound constantly satisfies.
For the second inequality, we define $f(c)=\frac{\binom{N}{A}-\binom{N-c}{A}}{c\cdot\binom{N}{A}}$, and we aim to show that $f(c)$ is monotonically decreasing.
Toward the goal, we reduce and prove the following inequality holds:
$\frac{\binom{N}{A}-\binom{N-c}{A}}{c\cdot\binom{N}{A}}\leq\frac{\binom{N}{A}-\binom{N-(c-1)}{A}}{(c-1)\cdot\binom{N}{A}}$ (19)
[…]
Next, we show the establishment of Equation (23) by mathematical induction:
$\bullet$ When $c=1$, we have $LHS=1+\frac{A}{N-A}$, while $RHS=1+\frac{A}{N-A}$. Apparently, the inequality holds.
$\bullet$ Let us assume that, when $c=k$, the conclusion holds, i.e., $1+\frac{kA}{N-k-A+1}\leq\prod_{i=0}^{k-1}\left(1+\frac{A}{N-A-i}\right)$. When $c=k+1$, we have
$\prod_{i=0}^{k}\left(1+\frac{A}{N-A-i}\right)$ $=\prod_{i=0}^{k-1}\left(1+\frac{A}{N-A-i}\right)\cdot\left(1+\frac{A}{N-A-k}\right)$ (24)
$\geq\left(1+\frac{kA}{N-k-A+1}\right)\cdot\left(1+\frac{A}{N-A-k}\right)$ (25)
$\geq 1+\frac{(k+1)A}{N-k-A}$ (26)
This shows that the inequality (i.e., Equation (23)) still holds, indicating that $f(c)$ is monotonically decreasing. As $c$ is an integer of $\{1,\dots,N-A\}$, we have
$f(c)\leq f(1)=\frac{\binom{N}{A}-\binom{N-1}{A}}{\binom{N}{A}}=\frac{A}{N}$ (27)
Apparently, we can attain the desired results as mentioned in Equation (18). ∎
@@ C.2 Preliminary
Suppose that Assumption 5.1, 5.2 and 5.3 hold. Let the local learning rate satisfy $\eta_{l}\leq\frac{1}{2L\sqrt{K(K+1)}}$. With FIARSE, we have the following conclusion:
@@
First, we establish a recursive relationship between $\mathbb{E}\left\|\boldsymbol{x}_{t,k+1}^{(i)}-\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}$ and $\mathbb{E}\left\|\boldsymbol{x}_{t,k}^{(i)}-\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}$:
@@
where the second term in Equation (30) is expanded to four parts: Equation (32), Equation (33), Equation (34), and Equation (35). Furthermore, applying Assumption 5.1 to Equation (32) and Assumption 5.3 to Equation (33), along with the equalities $\nabla_{\tilde{\boldsymbol{x}}_{t}}F_{i}\left(\tilde{\boldsymbol{x}}_{t}\right)\odot\mathcal{M}^{(i)}_{t}\left(\tilde{\boldsymbol{x}}_{t}\right)=\nabla F_{i}^{[0:\gamma_{i}]}\left(\tilde{\boldsymbol{x}}_{t}\right)$ and $\nabla F\left(\tilde{\boldsymbol{x}}_{t}\right)\odot\mathcal{M}^{(i)}_{t}\left(\tilde{\boldsymbol{x}}_{t}\right)=\nabla F^{[0:\gamma_{i}]}\left(\tilde{\boldsymbol{x}}_{t}\right)$, we simplify the above inequality for
$\mathbb{E}\left\|\boldsymbol{x}_{t,k+1}^{(i)}-\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}$ (36)
$\leq$ $\left(1+\frac{1}{K}+4(1+K)\eta_{l}^{2}L^{2}\right)\cdot\mathbb{E}\left\|\boldsymbol{x}_{t,k}^{(i)}-\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}$ (37)
$+4(1+K)\eta_{l}^{2}\left(\delta_{t}^{2}\left\|\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}+\mathbb{E}\left\|\nabla F_{i}^{[0:\gamma_{i}]}\left(\tilde{\boldsymbol{x}}_{t}\right)-\nabla F^{[0:\gamma_{i}]}\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}+\mathbb{E}\left\|\nabla F^{[0:\gamma_{i}]}\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}\right)$ (38)
Let $C$ be
$C=\delta_{t}^{2}\left\|\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}+\mathbb{E}\left\|\nabla F_{i}^{[0:\gamma_{i}]}\left(\tilde{\boldsymbol{x}}_{t}\right)-\nabla F^{[0:\gamma_{i}]}\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}+\mathbb{E}\left\|\nabla F^{[0:\gamma_{i}]}\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}.$ (39)
As $\eta_{l}\leq\frac{1}{2L\sqrt{K(K+1)}}$, we have
@@
Equation (42) holds because $\left(1+\frac{2}{K}\right)^{k}\leq 9$ for all $k\in\mathbb{N}_{+}$. Therefore, for all $k\in[K]$, we have
@@
With this conclusion, we can bound the left-hand side of Equation (28), which is
@@
Finally, by applying Assumption 5.2 to Equation (46), we achieve the desired conclusion. ∎
@@
Suppose that Assumption 5.1, 5.2 and 5.3 hold. Let the local learning rate satisfy $\eta_{l}\leq\min\left(\frac{1}{2L\sqrt{K(K+1)}},\frac{1}{6L\sqrt{(K+1)A}}\right)$. For any $t\in[T]$, $j\in\{0,\dots,n\}$, let $\boldsymbol{v}_{t}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}$ be
$\boldsymbol{v}_{t}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}=\nabla F^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right)-\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|K}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right)$ (47)
With FIARSE, we have the following conclusion:
$\sum_{j=0}^{n}\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\cdot\mathbb{E}\left\|\boldsymbol{v}_{t}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\right\|_{2}^{2}$ $\leq 4A\delta_{t}^{2}\left\|\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}+72KA\eta_{l}^{2}L^{2}\left(\sigma^{2}+\left\|\nabla F\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}\right)$ (48)
@@
In the beginning, let us define
@@
$\hat{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t}$ $=\nabla F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right)-\nabla_{\tilde{\boldsymbol{x}}_{t}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\odot\mathcal{M}^{(i)}_{t}\left(\tilde{\boldsymbol{x}}_{t}\right)\right),$ (49)
@@
$\tilde{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t,k}$ $=\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right)-\nabla_{\tilde{\boldsymbol{x}}_{t}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\odot\mathcal{M}^{(i)}_{t}\left(\tilde{\boldsymbol{x}}_{t}\right)\right).$ (50)
According to Equation (47), $\boldsymbol{v}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}$ is equivalent to
$\boldsymbol{v}_{t}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}=\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|}\sum_{i\in N_{\gamma_{j}^{\prime}}}\hat{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t}-\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|K}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\tilde{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t,k}$ (51)
To find a bound for Equation (48), we try to bound $\mathbb{E}\left\|\boldsymbol{v}\right\|_{2}^{2}$ using generalized mean inequality and Cauchy–Shwarz inequality: